When an electric motor is turned on at t = 0, its angular acceleration is š¯›¼ = 10e-0.5t where t is the time in seconds. What is the terminal angular velocity of the motor in rpm?


1
Expert's answer
2021-07-14T05:48:38-0400

"\alpha = 10 e^{-0.5t}= 10e^{-\frac{t}{2}}\\
\alpha = \frac{dw}{dt} =10e^{-\frac{t}{2}}\\
\int\alpha = \int\frac{dw}{dt} =\int10e^{-\frac{t}{2}}\\
\int _0^w dw =\int_{t=0} ^{t=\infin } 10e^{-\frac{t}{2}} dt\\
w =\frac{10e^{-\frac{t}{2}}}{-\frac{t}{2}} dt\\
w=-20e^{-\frac{t}{2}}|_0^{\infin}\\
w=20 rad/sec =190.985932 RPM"